# “Super-resolution” holographic optical tweezers array

### v1

@@ Abstract
Aligning light spots into arbitrary shapes is a fundamental challenge in holography, leading to various applications across diverse fields in science and engineering. However, as the spot interval approaches the wavelength of light, interference effects among the spots become prominent, which complicates the generation of a distortion-free alignment. Herein, we introduce a hologram design method based on the optimisation of a nonlinear cost function using a holographic phase pattern as the optimisation parameter. We confirmed a spot interval of 0.952(1) µm in a $5\times 5$ multispot pattern on the focal plane of a high-numerical-aperture (0.75) objective by observing the near-infrared (wavelength: 820 nm) holographic output light from a spatial light modulator device, a result which overcomes the limitation of a few micrometres under similar conditions. Furthermore, the definition of the Rayleigh diffraction limit is refined by considering the separation of spots and the spot interval, thereby concluding the achievement of “super-resolution.” The proposed method is expected to advance laser fabrication, scanning laser microscopy, and cold atom physics, among other fields.
Introduction
@@
Holography, which involves the control and utilisation of both the amplitude and phase of light, has provided numerous benefits—ranging from academic to practical applications—to the modern society for more than seventy years [1, 2]. The technological advancement of optical devices has accelerated the progress of holography; notably, the spatial light modulator (SLM) allows for the direct control of light wavefronts through the optical properties of liquid–crystal molecules [3]. Furthermore, the concept of a computer-generated hologram (CGH) facilitates the on-demand generation of desired light patterns by computing a phase pattern that reproduces the required output [4]. Holographic methods that modulate the light phase using SLMs are widely applied because of their high light utilisation efficiency [5, 6].
@@
Generating light spot arrays is another application of holography, with uses in laser fabrication [7, 8], scanning laser microscopy [9], and cold atom physics. Holographic optical tweezers are vital tools for aligning atoms in an array of arbitrary shapes [10, 11, 12, 13, 14], where the quality and spacing of the spot arrays affect the reproducibility of the atom alignment and the strength of interactions between atoms. Generating short-spaced array patterns without distortions is a technical challenge; however, this endeavour addresses a general and fundamental issue in optics, namely, the realisation of super-resolution [15, 16, 17, 18].
A common indicator used to determine whether adjacent light spots can be visually distinguished in microscopy is Rayleigh’s criterion [19]. This criterion states that two spots cannot be distinguished if their spacing is less than the distance from the centre to the first minimum of the “Airy disk,” which is the diffraction pattern of a circular aperture under uniform illumination. However, Rayleigh’s criterion is, at most, semiquantitative and applies only to the observation of incoherent point sources. In fact, under Rayleigh’s criterion, the two Airy disks appear almost merged, which raises the question of whether each spot functions appropriately as an independent optical tweezer. This directs us to reconsider or modify the criteria according to the context of the application.
@@
In this study, we propose a computer-generated hologram (CGH) design method that effectively generates a well-separated fine multispot light array. The proposed method is based on the optimisation of a nonlinear cost function, with the holographic phase pattern serving as the optimisation parameter. The cost function is defined as the correlation between a given target pattern and the output light pattern, which results from the light propagation from the phase distribution in the CGH. Additionally, we establish a quantitative super-resolution criterion for a multispot light array, based on the interrelationship between two conditions: spot spacing and separation of spots. The spot array is considered to achieve super-resolution if either the observed spot separation or the spacing is superior to that of two adjacent Airy disks aligned to satisfy the other conditions. By reproducing CGHs for a $5\times 5$ multispot pattern with near-infrared light ($\lambda=820$ nm), a spot spacing of $0.952(1)$ µm was confirmed for the focal plane of a high-numerical-aperture (NA$=0.75$) objective, which overcomes the limitation to a few micrometres under similar conditions.
Hologram design for fine spot arrays
@@
Figure 1: Hologram design concept. a, Schematic of CGH design procedure. On the SLM plane, user-defined phase $\theta_{j}$ is attached to incident light amplitude $a_{j}$ to make output light amplitude of SLM to be $\psi_{j}=a_{j}e^{i\theta_{j}}$. Light propagation, expressed symbolically by an illustration of a lens, acts as a unitary transformation on $\psi_{j}$ to yield light amplitude on the image plane, $\Psi_{i}$. Comparison of output pattern, $I_{i}=|\Psi_{i}|^{2}$, and target pattern, $T_{i}$, provides a guide to update $\theta_{j}$. b, Holographic phase patterns for the 5$\times$5 square lattice pattern optimised by CGM using FOI (left) and RSS (right) cost function.
@@
Figure 1a illustrates the concept of CGH design using a reflection-type SLM. For simplicity, the two-dimensional position on the SLM plane is denoted by $j$, which can be regarded as a vector index. The SLM applies a phase shift $\theta_{j}$ to the incident light amplitude $a_{j}$ at the $j$th position. Consequently, the output light amplitude from the SLM is expressed as $\psi_{j}=a_{j}\exp(i\theta_{j})$.
@@
The position in the image plane can be specified by introducing another vector $i$ to approximately represent the light propagating from the SLM to the image plane using the Fourier transform matrix $U=\{u_{ij}\}$. Therefore, the light amplitude on the image plane $\Psi_{i}$ is related to $\psi_{j}$ as follows:
@@
There are two methods for CGH design, which involve calculating the phase distribution $\{\theta_{j}\}$ to obtain the desired output patterns: self-consistent iterations and optimisation-based methods. A well-known example of the former is the iterative Fourier transform algorithm (IFTA) [6], which determines $\{\theta_{j}\}$ self-consistently by repeatedly performing light propagation calculations while imposing constraints on the light amplitudes on the SLM and image planes (Fig. 1a). In contrast, optimisation-based methods determine the optimal CGH according to the degree of correspondence between $I_{i}(\{\theta_{j}\})$ and $T_{i}$. Although optimisation-based methods require more computational resources, they offer higher computational stability and flexibility in designing CGHs.
The optimisation-based CGH design method utilising the analytical gradient of the cost function was first proposed by Harte [20]. In this approach, the set $\{\theta_{j}\}$ is determined by optimising the following cost function:
@@
which involves minimising the residual sum of squares (RSS) between $I_{i}$ and $T_{i}$. Note that $I_{i}$ and $T_{i}$ are normalised such that their sums equal unity, and $I_{i}$ is an explicit function of the phase distribution. This method provides holographic phase patterns similar to those obtained by IFTA. However, unique functionalities can be introduced into CGH design by selecting different cost functions [21]. In this study, we propose using the “fidelity of intensity” (FOI) as the cost function:
@@
where $\tilde{}$ denotes a quantity normalised with respect to its Euclidean norm.
Gradient-based optimisation methods including the conjugate gradient method (CGM) require computing the gradients of the cost function, which often becomes challenging for more complex functions than those considered in this study. Fortunately, $\partial f_{\text{FOI}}/\partial\theta_{j}$ can be explicitly expressed as follows:
@@
where $\mathcal{N}_{I}(\equiv\sqrt{\sum_{i}I_{i}^{2}})$ represents the Euclidean norm of $\{I_{i}\}$. This expression enables efficient coding that completes large-scale optimisation in a reasonable time, typically 10–20 min, starting with an arbitrary random pattern. Figure 1b illustrates the examples of designed CGHs obtained using the FOI and RSS methods, respectively. Despite the scepticism regarding the practical effectiveness of the FOI-designed CGH (Fig. 1b left), we assure that it generally performs well. See Methods for details (of the hologram design).
Figure 2: Hologram reproduction experiments and results. a. Schematic of experimental setup for hologram reproduction. Output from fibre-coupled laser diode (LD) is collimated by an achromatic lens ($\text{L}_{1}$) and is projected through a neutral density filter (ND) and an aperture (AP) on the active area of a spatial light modulator device (SLM), on which phase distribution of a displayed holographic phase pattern is attached. A $4f$-telescope system comprising a pair of achromatic lenses ($\text{L}_{2}$ and $\text{L}_{3}$) is utilised to project the SLM plane onto the entrance pupil of an objective lens ($\text{OB}_{1}$). The reproduced images are captured on the focal plane of $\text{OB}_{1}$, referred to as the “image plane,” using a CMOS image sensor (IS) through another objective lens of identical specifications ($\text{OB}_{2}$) and an achromatic lens ($\text{L}_{4}$) functioning as a tube lens. A pinhole (P) is placed at the focus of the $4f$ system to eliminate the zeroth-order and stray components. The numerical and observed images reproduced from FOI- and RSS-designed CGHs are presented as well. b, c: numerical results for b FOI- and c RSS-designed CGHs. d, e: experimental results of d FOI- and e RSS-designed CGHs. The spot spacing is set to 1.8 pixels in a common target pattern, corresponding to $0.952(1)$ µm on the image plane. All results were obtained without any adaptive modifications to the hologram patterns.
Results
Hologram reproduction and analyses. To investigate the properties of the proposed method, the reproduced images of RSS- and FOI-designed CGHs were compared for target patterns as $5\times 5$ square spot arrays with varying spot spacing. The reproduced images of CGHs were observed using the experimental setup illustrated in Fig. 2a. Figures 2b-2e display examples of numerical and experimental images reproduced from FOI- and RSS-designed CGHs for a target spot spacing of 1.8 pixels, which corresponds to 0.9225 µm in the experimental setup (refer to Methods).
When the spot spacing of the target pattern is less than two pixels, the spots in the numerically reproduced images begin to merge for the RSS-designed CGHs, whereas the results of FOI-designed ones maintain clear separation. However, in the experimental reproduction, even the images from FOI-designed CGHs cannot maintain spot separation for target spot spacing of less than 1.8 pixels.
Although adaptive adjustment of hologram patterns is frequently incorporated into CGH designs to enhance the image quality of holographic outputs [22, 23], the current results were obtained without any post-modification. In our experiments, a $5\times 5$ square spot array was selected as the target pattern to minimise the computational load, which increases with the number of spots, and to verify the principles of the proposed method.
The mean spot spacing is 0.952(1) µm for Fig. 2d, which cannot be attained by any other existing hologram design methods under similar optical conditions. Actually, reproduced patterns on the focal plane was observed by super-resolution optical pattern measurement [24, 25] with the consequence that the spot spacing was estimated as less than 1 µm using 852 nm trapping beam with a high-NA objective (NA=0.75 in design). Preliminary experiments on fluorescence image observation also confirmed trapping of atoms on the focal plane. These results will be reported elsewhere in the future.
@@ Results
Table 1: Properties of reproduced spot arrays. The spatial uniformity is assessed by the relative standard deviation (SD), which is the ratio of the SD to the mean value, in the powers of the $5\times 5$ light spots that form the spot array. The power of each light spot is estimated by applying fitting analysis to an output image (Methods). The light utilisation efficiency is calculated as the ratio of the output light power (measured by a power meter through a 1-mm-diameter pinhole placed on the image plane) to the illumination power measured on the SLM plane. The light utilisation efficiency is simply stated as the sum of $I_{i}$ over the area displayed in Fig. 2 for numerically reproduced images. The value and error of a physical quantity indicate the mean value and standard error (SE), respectively, in a series of quantities evaluated numerically or experimentally from $20$ CGHs, each designed by starting from a different initial random pattern to generate a common target.
Spot uniformity Utilisation efficiency

### v2

# Optimization-based hologram design for fine optical tweezers array and extension of super-resolution criteria

@@ Abstract
Aligning light spots into arbitrary shapes is a fundamental challenge in holography, leading to various applications across diverse fields in science and engineering. However, as the spot interval approaches the wavelength of light, interference effects among the spots become prominent, which complicates the generation of a distortion-free alignment. Herein, we introduce a hologram design method based on the optimization of a nonlinear cost function using a holographic phase pattern as the optimization parameter. We confirmed generation of a $5\times 5$ multispot pattern, in which each spot is separated clearly by a distance of 0.952(1) µm, on the focal plane of a high-numerical-aperture (0.75) objective by observing the near-infrared (wavelength: 820 nm) holographic output light from a spatial light modulator device. This result breaks through the spot spacing of a few micrometers that is a typical value achieved so far under similar experimental conditions. Furthermore, we propose redefinition of the Rayleigh diffraction limit by taking account of the separation of spot as well as the spot spacing. The proposed method is expected to advance laser fabrication, scanning laser microscopy, and cold atom physics, among other fields.
I Introduction
@@
Holography, which involves the control and utilization of both the amplitude and phase of light, has provided numerous benefits—ranging from academic to practical applications—to the modern society for more than seventy years [1, 2]. The technological advancement of optical devices has accelerated the progress of holography; notably, the spatial light modulator (SLM) allows for the direct control of light wavefronts through the optical properties of liquid-crystal molecules [3]. Furthermore, the concept of a computer-generated hologram (CGH) facilitates the on-demand generation of desired light patterns by computing a phase pattern that reproduces the required output [4]. Holographic methods that modulate the light phase using SLMs are widely applied because they enable dynamic pattern generation with high light utilization efficiency [5, 6].
@@
Among various applications of holography, generating light-spot arrays is one of the most popular examples for uses in laser fabrication [7, 8], scanning laser microscopy [9], and cold atom physics. Holographic optical tweezers are vital tools for aligning atoms in an array of arbitrary shapes [10, 11, 12, 13, 14], in particular, to construct “storage area” in neutral-atom quantum computers [15, 16]. Generating short-spaced, clearly-resolved light-spot arrays with CGHs is expected to benefit those applications; however, this is a technical challenge addressing a general and fundamental issue in optics, namely, the realization of super-resolution [17, 18, 19, 20].
CGHs for light-spot arrays are designed by a variety of methods, which are categorized into two main classes in terms of mathematical operations: self-consistent iterative methods and optimization-based methods. The former ones are fundamentally based on the Gerchberg-Saxton algorithm (GSA) [21], and in the field of hologram design, they are often referred to as iterative Fourier transform algorithms (IFTA) [6]. Popular CGH design methods such as weighted GSA (WGS) [22] and mixed-region amplitude-freedom (MRAF) [23] methods are regarded as the iterative methods with additional tricks. The optimization-based methods were proposed as ones advantageous in convergence and reproducibility, and an individual method is identified according to the choice of a cost function, a quantity to be optimized, and an optimization algorithm, such as direct search, simulated annealing, and Newtonian method. An optimal-rotation-angle (ORA) method [24] is classified as a variation of the optimization-based methods. All of the existing CGH design methods have own advantages but are, to a greater or lesser extent, ineffective for generating short-spaced light-spot arrays.
@@
In this study, we propose a CGH design method that effectively generates a well-resolved fine multispot light array. The proposed method is based on the optimization of a cost function, with the holographic phase pattern serving as the optimization parameter. The cost function is defined as the correlation between a given target pattern and the output light pattern, which results from the light propagation from the phase distribution in the CGH. This choice of a nonlinear cost function brings drastic improvement of holographic generation of short-spaced light-spot arrays compared to a similar method [25]. By reproducing CGHs for a $5\times 5$ multispot pattern with near-infrared light ($\lambda=820$ nm), a spot spacing of $0.952(1)$ µm was confirmed on the focal plane of a high-numerical-aperture ($\text{NA}=0.75$) objective, a result which shows notable reduction from a few micrometers typically observed under similar experimental conditions so far.
Additionally, we establish a quantitative super-resolution criterion for evaluating fineness of a multispot light array. Well-known resolution criteria, i.e., Rayleigh, Sparrow, and Abbe criteria, are established upon the spacing and separation of two adjascent light spot, however, the criteria are reduced to a boundary value only of the spot spacing. The proposed criterion is defined by the interrelationship between the spacing and separation, resultantly, to give a comprehensive one including those three criteria as well as importing the concept of modulation transfer function (MTF) [26] by emphasizing the role of image contrast (or separation) in super-resolution. Finally, the spot spacing obtained in this experiment does not satisfy the Rayleigh criterion, but analysis in conjunction with the spot separation showed that it satisfies our super-resolution criterion.
II Hologram design for fine spot arrays
@@
Figure 1: (a) Schematic of CGH design procedure. On the SLM plane, user-defined phase $\theta_{j}$ is attached to incident light amplitude $a_{j}$ to make output light amplitude of SLM to be $\psi_{j}=a_{j}e^{i\theta_{j}}$. Light propagation, expressed symbolically by an illustration of a lens, acts as a unitary transformation on $\psi_{j}$ to yield light amplitude on the image plane, $\Psi_{i}$. Comparison of output pattern, $I_{i}=|\Psi_{i}|^{2}$, and target pattern, $T_{i}$, provides a guide to update $\theta_{j}$. (b) CGHs for the $\times5$ square lattice pattern optimized by CGM using FOI (left) and RSS (right) cost function.
@@
Figure 1(a) illustrates the concept of CGH design using a reflection-type SLM. For simplicity, the two-dimensional position on the SLM plane is denoted by $j$, which can be regarded as a vector index. The SLM applies a phase shift $\theta_{j}$ to the incident light amplitude $a_{j}$ at the $j$th position. Consequently, the output light amplitude from the SLM is expressed as $\psi_{j}=a_{j}\exp(i\theta_{j})$.
@@
The position in the image plane can be specified by introducing another vector $i$ to approximately represent the light propagating from the SLM to the image plane which can generally be described by a unitary matrix $U=\{u_{ij}\}$, often referred to as the “transfer matrix.” This matrix is approximated by a Fourier transformation when the SLM and image planes are connected using imaging optics. Therefore, the light amplitude on the image plane $\Psi_{i}$ is related to $\psi_{j}$ as follows:
@@
As described in Introduction, we attempt to extend optimization-based methods that offer higher computational stability and flexibility in designing CGHs with a drawback of requirement of more computational resources. IFTA [6], the most popular example of the iterative methods, determines $\{\theta_{j}\}$ self-consistently by repeatedly performing light propagation calculations while imposing constraints on the light amplitudes on the SLM and image planes (Fig. 1(a)). In contrast, the optimization-based methods determine the optimal CGH according to the degree of correspondence between $I_{i}(\{\theta_{j}\})$ and $T_{i}$.
The optimization-based CGH design method utilizing the analytical gradient of the cost function was first proposed by Harte [25]. In this approach, the set $\{\theta_{j}\}$ is determined by optimizing the following cost function:
@@
which involves minimizing the residual sum of squares (RSS) between $I_{i}$ and $T_{i}$. Note that $I_{i}$ and $T_{i}$ are normalized as $\sum_{i}I_{i}=\sum_{i}T_{i}=1$, and $I_{i}$ is an explicit function of the phase distribution. As a merit of the optimization-based method, unique functionalities can be introduced into CGH design by selecting different cost functions. Bowman et al.[27] introduced a cost function based on a correlation between the target and output light amplitudes to achieve holographic complex-amplitude modulation. The final form of their cost function is a square of the light-amplitude correlation subtracted from unity, hence it can be regarded as a variation of RSS cost functions. It is also noted that their aim is simultaneous modulation of amplitude and phase, not the generation of fine patterns. Here, we propose using simply the correlation of intensity, i.e., “fidelity of intensity” (FOI):
@@
where $\tilde{}$ denotes a quantity normalized with respect to its Euclidean norm.
[…]
In the remaining part of this section, some remarks on numerical implementation of the proposed method are given. Among various numerical optimization algorithms, gradient methods become more advantageous in terms of computational complexity and convergence as the number of optimization parameters increases. In particular, the conjugate gradient method (CGM) is known as an efficient and stable one. While it is possible in principle to perform optimization by calculating gradients using numerical differentiation, we performed our calculations with analytical gradients because our design method becomes a large degree-of-freedom optimization problem involving a massive amount of computation. Fortunately, $\partial f_{\text{FOI}}/\partial\theta_{j}$ can be explicitly expressed as follows:
@@
where $\mathcal{N}_{I}(\equiv\sqrt{\sum_{i}I_{i}^{2}})$ represents the Euclidean norm of $\{I_{i}\}$. This expression enables efficient coding that completes large-scale optimization in a reasonable time, starting with an arbitrary random pattern.
Numerical optimization yielded holographic phase patterns as two-dimensional arrays of double-precision floating-point numbers within the interval $0,2\pi)$. As the SLM device is designed to handle 8-bit signals, the phase values were discretized into 256-step (8-bit) quantities for practical application. In this study, “CGHs” refers to the discretized versions. In a practical CGH design, the SLM plane is typically represented as an $N\times N$ square grid, where $N$ is a positive integer. The active area of the SLM is embedded within this grid by setting $a_{j}=0$ outside the active area. The image plane also becomes an $N\times N$ grid because the action of $U$ preserves the grid size. Suppose hologram reproduction is performed using light of wavelength $\lambda$ with imaging optics of effective focal length $f$. In that case, the actual size of the target can be estimated using the relation $\Delta x^{\prime}=\lambda f/(N\Delta x)$, which describes the relationship between the grid spacings on the SLM and image planes, $\Delta x$ and $\Delta x^{\prime}$, respectively.
III Experiments
Figure 2: Schematic of experimental setup for hologram reproduction. LD: laser diode, $\text{L}_{1}$-$\text{L}_{4}$: achromatic lens, ND: neutral density filter, AP: aperture, BS: beam splitter, P: pinhole, M: mirror, $\text{OB}_{1}$-$\text{OB}_{2}$: objective lens, IS: CMOS image sensor.
Experiments were conducted using the optical system depicted in Fig. 2. An 820 nm laser diode with a fiber port (LP820-SF80, Thorlabs) served as the light source. Output light from fiber port was collimated using an achromatic lens and was shaped through an aperture. The collimated light was directed onto a reflection-type liquid-crystal-on-silicon SLM (X13138-01 modified, Hamamatsu). Phase holograms [Fig. 1(b)] were displayed on the SLM surface that is filled up with $1272\times 1024$ pixels (pixel size: 12.5 µm $\times$ 12.5 µm). A blazed phase grating pattern was superimposed to demarcate the non-modulated (or, zeroth-order) light components from SLM outputs [28]. Since the blazed phase pattern was displayed inside a 360 px-radius circle around the center of the SLM surface, a radius of the effective aperture of the holographic outputs became 4.5 mm. Noting that the Gaussian radius of the illumination profile was estimated as approximately 20 mm on the SLM surface, the relative standard deviation of the illumination can be estimated as less than 0.030 over the effective aperture. Thus, the incident light profile can be regarded as a pseudo top-hat.
For generating holographic light-spot arrays, the SLM outputs were separated by a beam splitter and transferred through a 4ff telescope—0.8X-magnification optical system consisting of two achromatic lenses with a focal length of 125 mm and 100 mm—to the entrance pupil of an objective lens [MRH00401, Nikon: NA = 0.75 (air)]. If the illuminated area on the SLM plane (i.e., the SLM pixels labeled by $\{\forall j|a_{j}\neq 0\}$) is not incident on the entrance pupil of the objective lens, it reduces the effective NA and fails to generate a short spot spacing, thereby limiting the performance of the objective. Therefore, a suitable choice of the illumination area and/or magnification of the transfer optical systems was required. Focal images were observed using a 60X-magnification optical system, comprising the same objective lens used for focusing, an achromatic lens with a focal length of 300 mm, and an image sensor (Lumenera, LU135M-IO-WOIR).
We designed CGHs by applying the CGM to minimize the FOI cost function for given target patterns. The number of CGM iterations was set at 1000, which is typically sufficient for the relative decrease in FOI cost per iteration step to be less than $10^{-8}$. The minimization-based hologram design problem does not necessarily yield unique solutions. Therefore, the designed CGHs depend on the initial patterns selected, which are prepared by assigning a pseudorandom number (in the range $0,2\pi))$ to each lattice point on the grid in the SLM plane.
Here, we assume that the hologram and the target planes are connected using Fourier optics. The input and target planes are represented as spatial grids of the same size in the numerical Fourier transform, with the target plane comprising $\times800$ points. To increase the effective resolution of the target image by a factor of ten, the method described in [29] was applied to the SLM plane. Specifically, the hologram region is embedded into the center of an $\times8000$ grid, where the light amplitude is set to zero outside the hologram. Moreover, we assume that the illumination is a top-hat laser beam, indicating that $a_{j}$ in Fig. 1(a) presumes a nonzero constant value within the central circular region of a 360-pixel radius.
Ultimately, the hologram design becomes a 640,000-dimensional numerical optimization problem. Most computing resources are consumed by repeated evaluations of the cost function and its derivatives, which involve the propagation calculations of the two-dimensional light amplitude distribution expressed by $\times8000$ components. To efficiently perform such large-scale propagation computations, we applied a double-precision complex FFT algorithm running on general-purpose graphics processing units, achieving reasonable computational times (typically 10–20 minutes).
Figure 1(b) illustrates examples of the designed CGHs obtained using the FOI and RSS methods, respectively. The CGH designed with the FOI method appears to exhibit an irregular and rough phase front to raise a question whether it is merely a mathematical object without physical meaning. Nevertheless, as will be comfirmed later, the FOI method is superior to the RSS method in terms of both uniformity and separability.
Figure 3: (a), (b): Numerically reproduced images of square lattice target patterns with different spot spacings. The upper row was calculated from FOI-designed CGHs, and the lower row from RSS-designed ones. The spot spacings are 1.6, 1.8, 2.1, and 2.5 pixels from left to right. (c), (d): Experimental results of images reproduced from FOI- and RSS-designed CGHs, respectively. The spot spacing was set to 1.8 pixels in a common target pattern, corresponding to an actual length of $0.952(1)$ µm on the image plane.
@@ III Experiments
Table 1: Properties of reproduced spot arrays. The spatial uniformity is assessed by the relative standard deviation (SD), which is the ratio of the SD to the mean value, in the powers of the $5\times 5$ light spots that form the spot array. The light utilization efficiency is calculated as the ratio of the output light power (measured by a power meter through a 1-$m$-diameter pinhole placed on the image plane) to the illumination power measured on the SLM plane. The light utilization efficiency is simply stated as the sum of $I_{i}$ over the area displayed in Fig. 3 for numerically reproduced images. The value and error of a physical quantity indicate the mean value and standard error (SE), respectively, in a series of quantities evaluated numerically or experimentally from 2020 CGHs, each designed by starting from a different initial random pattern to generate a common target. For numerical calculations with aberrations, see the main texts in Sec. IV.2.
Spot uniformity Utilization efficiency
@@
IV Results
IV.1 Hologram reproduction and analyses
To investigate the properties of the proposed method, the reproduced images of RSS- and FOI-designed CGHs were compared for target patterns of $5\times 5$ square spot arrays. Figures 3(a) and (b) respectively show numerically reproduced images of the RSS- and FOI-designed CGHs for target patterns with spot spacings of 1.6, 1.8, 2.1, and 2.5 pixels. When the spot spacing of the target pattern becomes less than 2.1 pixels, the spots in the numerically reproduced images begin to merge for the RSS-designed CGHs, whereas the results of FOI-designed ones maintain clear separation. However, even in the numerical reproduction, the spot separation breaks down in the reproduced images of the FOI-designed CGHs for target spot spacing of less than 1.8 pixels.
The output images reproduced from CGHs were observed using the experimental setup illustrated in Fig. 2. Spot spacing of target patterns are set to 1.8 pixels, which corresponds to 0.9225 µm in the image plane, for observing a distinct difference between results of the FOI and RSS methods. We generated 20 holograms by applying both RSS and FOI methods starting with different random patterns as initial conditions in optimization calculation. Figures 3(c) and (d) display examples of reproduced images of the FOI- and RSS-designed holograms, respectively. We also tried to reproduce these images for target spot-interval of less than 1.8 pixels experimentally, but are noticed that it would be difficult to obtain well-resolved multispot patterns.
The spacing and positions of the spots were determined by applying fitting analysis to the output patterns on the image plane. We adopt the following model function for an incoherent superposition of Gaussian spots with a common spot radius, $w_{0}$, on uniform background $B$:
$I_{\text{fit}}=\sum_{k}A_{k}\exp\left[-\frac{(x-x_{k})^{2}+(y-y_{k})^{2}}{w_{0}^{2}/2}\right]+B,$ (5)
where $k$ denotes an index to distinguish the spots, and the peak “height” and position of the $k$-th spot is expressed as $A_{k}$ and $(x_{k},y_{k})$.
The tilt angle of an image, $\theta$, is included in the practical fitting calculations. Consequently, the fitting model becomes a (three times the number of spots plus three)-dimensional problem, rendering the analysis as computationally demanding. Therefore, the quantitative analyses were restricted to a $5\times 5$ square spot alignment with spacing $d$. After positioning a peak at $(X_{0},Y_{0})$, the other $(x_{k},y_{k})$ positions are automatically determined as $(X_{0}+md,Y_{0}+nd)$ for pairs of suitable integers mm and $n$, depending on the geometrical configuration of the target spot array.
The number of trials in the analyses of RSS results was reduced because one of the CGHs obtained using the RSS method was excluded from the experiments and analyses due to the failure of convergence in the hologram design. Although the RSS method successfully yields CGHs, the reproduced images of the RSS-designed holograms exhibit significant distortions, as depicted in Figs. 3(b) and 3(d). In such cases, the aforementioned analysis often fails to determine the peak positions in the reproduced images. Therefore, we adopted the results of an FOI-designed hologram that was obtained using the same initial random pattern.
The mean spot spacing is 0.952(1) µm for Fig. 3(c), which cannot be attained by any other existing hologram design methods under similar optical conditions. Actually, reproduced patterns on the focal plane was observed by super-resolution optical pattern measurement [30, 31] with the consequence that the spot spacing was estimated as less than 1 µm using 852 nm trapping beam with a high-NA objective (NA=0.75 in design). Although adaptive adjustment of CGHs is frequently incorporated into CGH designs to enhance the image quality of holographic outputs [32, 33], the current results were obtained without any post-modification.
Preliminary experiments on fluorescence image observation also confirmed trapping of atoms on the focal plane. These results will be reported elsewhere in the future.
IV.2 Spot uniformity and light utilization efficiency